# A model-agnostic gravitational-wave background characterization algorithm

### v1

@@ Abstract
As ground-based gravitational-wave (GW) detectors improve in sensitivity, gravitational-wave background (GWB) signals will progressively become detectable. Currently, searches for the GWB model the signal as a power law, however deviations from this model will be relevant at increased sensitivity. Therefore, to prepare for the range of potentially detectable GWB signals, we propose an interpolation model implemented through a transdimensional reverse-jump Markov chain Monte Carlo (RJMCMC) algorithm. This interpolation model foregoes a specific physics-informed model (of which there are a great many) in favor of a flexible model that can accurately recover a broad range of potential signals. In this paper, we employ this framework for an array of GWB applications. We present three dimensionless fractional GW energy density injections and recoveries as examples of the capabilities of this spline interpolation model. We further demonstrate how our model can be implemented for hierarchical GW analysis on $\Omega_{\textrm{GW}}$.
@@ I Introduction
Gravitational-wave (GW) detection is rapidly evolving into a precision science involving a variety of measurement techniques and detection strategies. In contrast to the individual compact binary signals observed with ground-based GW detectors thus far [1], the gravitational-wave background (GWB) is the superposition of all subthreshold signals that cannot be individually resolved by the detector network. Recently, multiple pulsar timing array (PTA) collaborations presented evidence for a GWB in the nHz frequency band [2, 3, 4, 5]. That signal is consistent with the predictions for a population of supermassive black hole binaries [6]. No gravitational wave background has been detected in the 10-1000 Hz band, where detectors such as Advanced LIGO [7], Advanced Virgo [8], and KAGRA operate [9]; however, such a discovery could provide a wealth of information.
@@
There is no shortage of predicted GWB signals in the ground-based detector band. On the astrophysical side, a signal from unresolved compact binary coalescences (CBCs) carries information about populations of binary black holes (BBHs) [10], high redshift metallicity [11], and populations of black holes in globular clusters [12, 13]. On the cosmological side, signals from first-order phase transitions [14, 15, 16, 17, 18, 19, 20, 21, 22], cosmic strings [23, 24, 25, 26], or inflationary models [27], among many others, have also been proposed [28]. Generally, the CBC signal is expected to dominate the band, and predictions show that detection is possible with improvements to the current detector network [29].
@@
The canonical model for the GWB spectrum from 10-1000 Hz is a power law (PL), and there are two primary reasons for adopting this model. Firstly, current interferometers (often referred to as second generation – 2G – detectors) are not sufficiently sensitive to GWs  [^1] Currently only very loud and/or nearby individual signals can be detected, which may be viewed as outliers of the total GW emitter population. The sensitivity to subthreshold background signals is therefore low; this is further compounded by the fact that the GWB is usually targeted using a search for persistent isotropic background signals [29], which is systematically less sensitive to a CBC background in this frequency band, as this signal is neither persistent [30] nor isotropic [31], so an initial detection of the GWB will likely be possible using a simple model that reflects that sensitivity. Secondly, most models for a GWB predict a (approximate) PL shape in the ground-based detector band. For the unresolved CBC background, for example, it is predicted that most of the GW power in the sensitive frequency band comes from the early inspiral phase of compact binaries, with a gradual turnover in the spectrum at higher frequencies, where the detectors are less sensitive. However, there are several reasons to expect deviations from this typical PL assumption: these include sensitivity improvements in upcoming detectors [32, 33, 34], deviations due to specific population models or waveform choices [35], distinctive features in cosmological stochastic signals [14, 15, 16, 17, 18, 19, 20, 21, 22, 36], and the possibility of distinguishing multiple correlated signals [37, 38, 39].
@@
The detection and characterization of the GWB spectrum can be leveraged to infer properties of the stellar-mass black hole population. This includes, for example, probing the merger rate as a function of cosmological redshift [40, 35], typically using a Madau-Dickinson-like merger rate evolution [10, 29]. One expects deviations from this fiducial model through potential sub-populations of primordial black holes as well as uncertainties in the star formation rate at high redshifts (see Sec. II.2).
@@
In this paper, we outline the scientific case for going beyond the PL assumption for GWBs and the Madau-Dickinson-like redshift evolution, and we introduce a framework for detecting and characterizing the GWB or the BBH merger rate evolution based on the reversible jump Markov chain Monte Carlo (RJMCMC) algorithm [41]. For modeling the GWB spectrum, for example, we interpolate a set of “knots” arranged in frequency space to the frequency bins of the measured data. The amplitude of the GWB and the location in the frequency space of each knot, as well as the number of knots, are independent parameters in the model. By making the number of knots variable, the interpolation model can fit a variety of spectral shapes and can yield a detection statistic for a GWB. For the merger rate as a function of redshift, we use the same flexible approach, instead placing the knots in redshift space.
@@
We focus on ground-based detector spectral ranges and use Advanced LIGO, Advanced Virgo, and Cosmic Explorer [42, 43] design sensitivities in our data analysis examples. The methods can be applied to other frequency bands and detection techniques, and will have valuable applications for both PTAs and the upcoming Laser Interferometer Space Antenna (LISA).
@@
The rest of this paper is organized as follows. In Sec. II we discuss current models for the GWB, including astrophysical models, and make the case for using a more flexible model. In Sec. III we review search methods for a GW background, while in Sec. IV we discuss the spline interpolation model for the GWB and the transdimensional fitting algorithm. In Sec. V we apply our new spectral models to two realistic examples for the CBC-generated GWB, and to a generic BPL background. We show results on simulated data for both current and next-generation detectors. In Sec. VI we demonstrate how to use the new framework to fit the BBH merger rate as a function of redshift using only measurements from the GWB searches. In Sec. VII we discuss current and future applications of this method and offer concluding remarks.
@@ II The GWB signal
where $\rho_{\rm gw}$ is the energy density in gravitational waves in the universe, $\rho_{c}=3c^{2}H_{0}^{2}/8\pi G$ is the critical energy density to close the universe and $H_{0}$ is the Hubble rate today. We use $H_{0}=67.66\,\mathrm{km/s/Mpc}$ from the Planck satellite 2018 results [44]. In some papers the astrophysical GWB is characterized in terms of the GW flux (e.g. [45, 46]), but we choose to express the signal in terms of $\Omega_{\textrm{GW}}$ here for comparison with past results from ground-based experiments.
@@ II.1 Current models for the GWB Spectrum
[…]
$\Omega_{\textrm{GW}}(f)=\Omega_{\textrm{ref}}\left(\frac{f}{f_{\mathrm{ref}}}\right)^{\alpha},$ (2)
with spectral index $\alpha=2/3$ and reference amplitude $\Omega_{\mathrm{ref}}$, which depends upon the population of systems that produce the signal, and is typically left as a free parameter. This prediction is based on waveforms computed using a zeroth-order post-Newtonian expansion [47], under the assumption that each binary is far from merger.
Other GWB models, such as those arising from slow-roll inflation [28] or cosmic strings [23, 24, 25, 26], also predict PL spectra, but each with a different spectral index. To accommodate the range of potential PL backgrounds, the joint posterior distribution $p[\Omega_{\mathrm{ref}},\alpha|\hat{C}(f)]$ is usually fit using Markov chain Monte Carlo (MCMC) methods, where $\hat{C}(f)$, described in detail in Sec. III, is the measured (normalized) cross-correlation spectrum between detectors. More complex models are also considered, including BPL spectra to approximate features from first-order phase transitions in the early Universe [14, 15, 16, 17, 18, 19, 20, 21, 22], and models that change the overlap reduction function due to the inclusion of non-tensorial polarizations [48, 49].
Recent computational advances have enabled population-based modeling of the GWB and the simultaneous fitting of individual events and the unresolved background [50, 40, 51, 29, 52, 53]. In this approach, the unresolved (or unsubtracted) CBC population is fit directly to the cross-correlation spectrum $\hat{C}(f)$, with parameterized priors on masses, spins, and redshifts. A “reference bank” of waveforms allows rapid adjustments to a fiducial spectral model based on changes in population assumptions [52, 35]. Although these models are powerful, their performance is highly dependent on prior choices and population assumptions, complicating their interpretation.
Using these techniques, one can fit the merger rate as a function of cosmological redshift [40, 35], typically assuming that binaries form in the field so the merger rate follows the star formation rate convolved with some time delay between formation and merger. This leads to a Madau-Dickinson-like merger rate evolution [10, 29]. One can also potentially constrain the metallicity dependence on the evolution of BBHs [52], and probe the contribution to the GWB from hierarchical mergers [13].
@@ II.2 The Case for a Flexible Model
When the background is dominated by CBCs, the frequency at which the spectrum turns over and the deviations from the PL shape depend on the distribution of masses of the merging binaries, along with the redshift at which they merge. Populations of compact binaries merging at different frequencies produce unique imprints in the overall background spectrum – for example, the GWB due to a population of heavy BBHs at high redshifts will have a turnover at lower frequencies than a population of binary neutron star (BNS) or neutron star-black hole (NSBH) mergers. These features can be probed directly in background searches. Furthermore, if these populations have roughly comparable contributions at lower frequencies (e.g. $\sim 10-100$ Hz), then the turnover in the spectrum due to one population could distort the shape of the combined spectrum from all sources, requiring joint modeling.
@@
Furthermore, recent studies that use more realistic waveforms to make GWB predictions challenge the simple PL assumption at low frequencies because the contribution of the late inspiral phase will change the shape of the spectrum – especially the specific value of the spectral index [35]. However, detecting such a deviation will likely require next generation detectors. In this context, as detectors become more sensitive and resolvable CBC signals are subtracted before searching for a GWB, the residual GWB after subtraction may no longer take the same expected PL shape [32, 33, 34].
@@
Looking beyond CBC backgrounds, models like first-order phase transitions in the early universe predict a BPL shape, where the spectral indexes before and after the break correspond to physically predictable quantities, and the frequency where the spectral index changes corresponds to the temperature of the universe when the phase transition took place [14, 15, 16, 17, 18, 19, 20, 21, 22, 36]. Finally, correlated magnetic noise between detectors could have a complicated spectral shape that will need to be characterized and separated from a true GWB signal [37, 38, 39].
@@
Next, we turn to the merger rate vs. redshift distribution for binary black holes that contribute a GW background. Deviations from the typical Madau-Dickinson-like model could come due to lensing of high redshift black holes [54, 55, 56], excess star formation at high redshifts hinted at in recent JWST results [57, 58], or the effects of metallicity on the formation of black holes at high redshift [59, 60, 61, 62, 63, 64, 52].
@@
All of these added complexities, which can potentially be probed with next generation detectors, motivate a more flexible model than a simple PL. In this paper, we use linear and spline interpolation with a variable number of knots to flexibly model the GWB spectrum and the BBH merger rate as a function of redshift. Although it may seem less informative compared to physical models used in parameter estimation, our model avoids the pitfalls of increased prior volume (and corresponding Occam penalties), reduces trials factors associated with testing multiple models, and could be used to fit correlated noise in the future as well.
@@ III Search methods overview
In this section we outline the main methods used to search for a GW background with ground-based detectors—readers already familiar with this material can skip to Sec. IV.
Common assumptions that simplify stochastic GW analyses are that the GWB is isotropic, broadband, stationary, Gaussian, and unpolarized [65]. In this limit, a sufficient search strategy is to cross-correlate data from pairs of instruments. Although there are search strategies for detecting and characterizing a GWB that is non-Gaussian [66, 50, 67, 30], anisotropic [68, 69, 70, 71, 72, 73, 74], narrowband [69], or some combination of each [75, 76, 77, 78, 79], we focus on the isotropic, Gaussian, unpolarized and stationary case.
A hybrid frequentist-Bayesian approach is used to perform the search [80, 81, 82, 83]. First, an unbiased frequentist estimator for the strength of the background in individual frequency bins is constructed by cross-correlating many chunks of data [65, 81, 83],
$\hat{C}_{ij}(f;t)=\frac{2}{T}\frac{\mathrm{Re}\left[\tilde{s}_{i}^{*}(f;t)\tilde{s}_{j}(f;t)\right]}{\Gamma_{ij}(f)S_{0}(f)},$ (3)
where $\tilde{s}_{i}(f;t)$ is the Fourier transform of the strain data in detector $i$, starting at time $t$ and ending at time $t+T$, with $T=192\,\mathrm{s}$ in most recent analyses [29]. $\Gamma_{ij}(f)$ is the overlap reduction function [84, 85], and $S_{0}(f)=3H_{0}^{2}/(10\pi^{2}f^{3})$ is a normalization factor that makes $\langle\hat{C}_{ij}(f)\rangle=\Omega_{\textrm{GW}}(f)$ (in the absence of correlated noise [37]).
In the limit where the GW strain amplitude is much smaller than the typical noise in the detectors (on the short, $\mathcal{O}(100\,\mathrm{s})$ timescales that $\hat{C}_{ij}(f;t)$ is calculated over), the variance of $\hat{C}_{ij}(f;t)$ is given by [81]:
@@
where $\Delta f$ is the width of the frequency bins used in the analysis (0.3125 Hz in most recent LVK collaboration results, [29]), and $P_{i}(f;t)$ is the one-sided power spectral density of the noise in detector $i$.
@@
In situations where the detector noise does not dominate the signal, $\sigma^{2}(f)$ (or potentially even $\sigma_{ij}^{2}(f;t)$) would need to be estimated either independently, or as a part of the Bayesian step we discuss below. This is likely to be the situation for third-generation detectors, where we may no longer be in the weak signal regime. In this work, we do not account for the intermediate or large signal regime. We assume that we have an independent measurement of the detector noise PSDs. We discuss potential methods for handling the intermediate and strong signal regimes in Section VII.
Next, the frequentist statistic is used to perform a weighted average across frequency bins assuming a specific spectral model [65, 81, 83], to estimate the background signal at a reference frequency. The same spectral model that fits $\hat{C}(f)$ for the spectral shape and amplitude of the background can also be used in Bayesian parameter estimation [80, 48, 37, 83]. The $\hat{C}(f)$ spectrum is obtained by averaging many time segments together, and so the central limit theorem implies that it should be Gaussian distributed. Therefore, the likelihood used for a Bayesian analysis of $\hat{C}(f)$ is given by [80, 81, 82]
@@
While we discussed the broad outline of the search method in this section, for the injection studies in Sections V we do not simulate individual time segments or time domain data—instead we calculate $\sigma^{2}(f)$ using publicly available sensitivity curves, and simulate the final $\hat{C}(f)$ using this variance.

### v2

# Model-agnostic gravitational-wave background characterization algorithm

@@ Abstract
As ground-based gravitational-wave (GW) detectors improve in sensitivity, gravitational-wave background (GWB) signals will progressively become detectable. Currently, searches for the GWB model the signal as a power law; however, deviations from this model will be relevant at increased sensitivity. Therefore, to prepare for the range of potentially detectable GWB signals, we propose an interpolation model implemented through a transdimensional reversible-jump Markov chain Monte Carlo algorithm. This interpolation model foregoes a specific physics-informed model (of which there are a great many) in favor of a flexible model that can accurately recover a broad range of potential signals. In this paper, we employ this framework for an array of GWB applications. We present three dimensionless fractional GW energy density injections and recoveries as examples of the capabilities of this spline interpolation model. We further demonstrate how our model can be implemented for hierarchical GW analysis on $\Omega_{\textrm{GW}}$.
@@ I Introduction
Gravitational-wave (GW) detection is rapidly evolving into a precision science involving a variety of measurement techniques and detection strategies. In contrast to the individual compact binary signals observed with ground-based GW detectors thus far [Abbott_2023], the gravitational-wave background (GWB) is the superposition of all subthreshold signals that cannot be individually resolved by the detector network. Recently, multiple pulsar timing array (PTA) collaborations presented evidence for a GWB in the nanohertz frequency band [NANOGrav:2023gor, EPTA:2023fyk, PPTA, CPTA]. That signal is consistent with the predictions for a population of supermassive black hole binaries [agazieNANOGrav15Yr2023a]. No gravitational-wave background has been detected in the 10100010-1000 Hz band, where detectors such as Advanced LIGO [LIGOScientific:2014pky], Advanced Virgo [Acernese2015], and KAGRA operate [KAGRA_design_doc]; however, such a discovery could provide a wealth of information.
@@
There is no shortage of predicted GWB signals in the ground-based detector band. On the astrophysical side, a signal from unresolved compact binary coalescences (CBCs) carries information about populations of binary black holes (BBHs) [PhysRevX.13.011048], high redshift metallicity [fishbach2021], and populations of black holes in globular clusters [Romero-Shaw:2020siz, Kou:2024gvp]. On the cosmological side, signals from first-order phase transitions [Grischuk1975, Hogan:1986dsh, Huber:2008hg, Child:2012qg, Hindmarsh:2013xza, Hindmarsh:2015qta, Hindmarsh:2017gnf, Jinno:2016vai, Weir:2017wfa], cosmic strings [Battye1997, vilenkinCosmicStringsOther2000, Jeannerot:2003qv, Damour:2004kw], or inflationary models [Guzzetti:2016mkm], among many others, have also been proposed [Caprini2018]. Generally, the CBC signal is expected to dominate the band, and predictions show that detection is possible with improvements to the current detector network [KAGRA:2021kbb].
@@
The canonical model for the GWB spectrum from 10100010-1000 Hz is a power law (PL), and there are two primary reasons for adopting this model. Firstly, current interferometers (often referred to as second-generation –2G– detectors) are not sufficiently sensitive to GWs,  [^1] Currently only very loud and/or nearby individual signals can be detected, which may be viewed as outliers of the total GW emitter population. The sensitivity to subthreshold background signals is therefore low; this is further compounded by the fact that the GWB is usually targeted using a search for persistent isotropic background signals [KAGRA:2021kbb], which is systematically less sensitive to a CBC background in this frequency band, as this signal is neither persistent [Lawrence:2023buo] nor isotropic [Jenkins_2018]. so an initial detection of the GWB will likely be possible using a simple model that reflects that sensitivity. Secondly, most models for a GWB predict a (approximate) PL shape in the ground-based detector band. For the unresolved CBC background, for example, it is predicted that most of the GW power in the sensitive frequency band comes from the early inspiral phase of compact binaries, with a gradual turnover in the spectrum at higher frequencies, where the detectors are less sensitive. However, there are several reasons to expect deviations from this typical PL assumption: these include sensitivity improvements in upcoming detectors [Cutler:2005qq, Sachdev:2020bkk, Sharma:2020btq], deviations due to specific population models or waveform choices [Renzini:2024pxt], distinctive features in cosmological stochastic signals [Grischuk1975, Hogan:1986dsh, Huber:2008hg, Child:2012qg, Hindmarsh:2013xza, Hindmarsh:2015qta, Hindmarsh:2017gnf, Jinno:2016vai, Weir:2017wfa, Martinovic:2020hru], and the possibility of distinguishing multiple correlated signals [Meyers:2020qrb, janssensImpactSchumannResonances2021, janssensCorrelated11000Hz2023].
@@
The detection and characterization of the GWB spectrum can be leveraged to infer properties of the stellar-mass black hole population. This includes, for example, probing the merger rate as a function of cosmological redshift [Callister:2020arv, Renzini:2024pxt], typically using a Madau-Dickinson-like merger rate evolution [PhysRevX.13.011048, KAGRA:2021kbb]. One expects deviations from this fiducial model through potential subpopulations of primordial black holes as well as uncertainties in the star formation rate at high redshifts (see Sec. II.2).
@@
In this paper, we outline the scientific case for going beyond the power law assumption for GWBs and the Madau-Dickinson-like redshift evolution, and we introduce a framework for detecting and characterizing the GWB or the BBH merger rate evolution based on the reversible jump Markov chain Monte Carlo (RJMCMC) algorithm [greenReversibleJumpMarkov1995]. For modeling the GWB spectrum, for example, we interpolate a set of “knots” arranged in frequency space to the frequency bins of the measured data. The amplitude of the GWB and the location in the frequency space of each knot, as well as the number of knots, are independent parameters in the model. By making the number of knots variable, the interpolation model can fit a variety of spectral shapes and can yield a detection statistic for a GWB. For the merger rate as a function of redshift, we use the same flexible approach, instead placing the knots in redshift space.
@@
We focus on ground-based detector spectral ranges and use Advanced LIGO, Advanced Virgo, and Cosmic Explorer [Evans:2021gyd, Evans:2023euw] design sensitivities in our data analysis examples. The methods can be applied to other frequency bands and detection techniques and will have valuable applications for both PTAs and the upcoming Laser Interferometer Space Antenna (LISA).
@@
The rest of this paper is organized as follows. In Sec. II we discuss current models for the GWB, including astrophysical models, and make the case for using a more flexible model. In Sec. III we review search methods for a GW background, while in Sec. IV we discuss the spline interpolation model for the GWB and the transdimensional fitting algorithm. In Sec. V we apply our new spectral models to two realistic examples for the CBC-generated GWB and to a generic broken power law (BPL) background. We show results on simulated data for both current and next-generation detectors. In Sec. VI we demonstrate how to use the new framework to fit the BBH merger rate as a function of redshift using only measurements from the GWB searches. In Sec. VII we discuss current and future applications of this method and offer concluding remarks.
@@ II The GWB signal
where $\rho_{\rm gw}$ is the energy density in gravitational waves in the Universe, $\rho_{c}=3c^{2}H_{0}^{2}/8\pi G$ is the critical energy density to close the Universe and $H_{0}$ is the Hubble rate today. We use $H_{0}=67.66$ km/s/Mpc from the Planck satellite 2018 results [Planck:2018vyg]. In some papers the astrophysical GWB is characterized in terms of the GW flux (e.g. [KAGRA:2021mth, Bellie:2023jlq]), but we choose to express the signal in terms of $\Omega_{\textrm{GW}}$ here for comparison with past results from ground-based experiments.
@@ II.1 Current models for the GWB spectrum
If the dominant contribution to the GWB comes from the early inspiral phase of CBCs, the spectrum is expected to be a PL in the frequency band $\mathcal{O}(10–100\,\textrm{Hz})$, where current ground-based detectors are most sensitive:
$\Omega_{\textrm{GW}}(f)=\Omega_{\textrm{ref}}\left(\frac{f}{f_{\textrm{ref}}}\right)^{\alpha},$ (2)
with spectral index $\alpha=2/3$ and reference amplitude $\Omega_{\textrm{ref}}$, which depends upon the population of systems that produce the signal, and is typically left as a free parameter. This prediction is based on waveforms computed using a zeroth-order post-Newtonian expansion [Ajith:2007kx], under the assumption that each binary is far from merger.
Other GWB models, such as those arising from slow-roll inflation [Caprini2018] or cosmic strings [Battye1997, vilenkinCosmicStringsOther2000, Jeannerot:2003qv, Damour:2004kw], also predict PL spectra, but each with a different spectral index. To accommodate the range of potential PL backgrounds, the joint posterior distribution $p[\Omega_{\textrm{ref}},\alpha|\hat{C}(f)]$ is usually fit using Markov chain Monte Carlo (MCMC) methods, where $\hat{C}(f)$, described in detail in Sec. III, is the measured (normalized) cross-correlation spectrum between detectors. More complex models are also considered, including BPL spectra to approximate features from first-order phase transitions in the early Universe [Grischuk1975, Hogan:1986dsh, Huber:2008hg, Child:2012qg, Hindmarsh:2013xza, Hindmarsh:2015qta, Hindmarsh:2017gnf, Jinno:2016vai, Weir:2017wfa], and models that change the overlap reduction function due to the inclusion of nontensorial polarizations [Callister:2017ocg, Tsukada:2023stv].
Recent computational advances have enabled population-based modeling of the GWB and the simultaneous fitting of individual events and the unresolved background [smithOptimalSearchAstrophysical2018, Callister:2020arv, KAGRA:2021duu, KAGRA:2021kbb, Turbang:2023tjk, Lalleman:2025xcs]. In this approach, the unresolved (or unsubtracted) CBC population is fit directly to the cross-correlation spectrum $\hat{C}(f)$, with parametrized priors on masses, spins, and redshifts. A “reference bank” of waveforms allows rapid adjustments to a fiducial spectral model based on changes in population assumptions [Turbang:2023tjk, Renzini:2024pxt]. Although these models are powerful, their performance is highly dependent on prior choices and population assumptions, complicating their interpretation.
Using these techniques, one can fit the merger rate as a function of cosmological redshift [Callister:2020arv, Renzini:2024pxt], typically assuming that binaries form in the field so the merger rate follows the star formation rate convolved with some time delay between formation and merger. This leads to a Madau-Dickinson-like merger rate evolution [PhysRevX.13.011048, KAGRA:2021kbb]. One can also potentially constrain the metallicity dependence on the evolution of BBHs [Turbang:2023tjk], and probe the contribution to the GWB from hierarchical mergers [Kou:2024gvp].
@@ II.2 The case for a flexible model
When the background is dominated by CBCs, the frequency at which the spectrum turns over and the deviations from the PL shape depend on the distribution of masses of the merging binaries, along with the redshift at which they merge. Populations of compact binaries merging at different frequencies produce unique imprints in the overall background spectrum – for example, the GWB due to a population of heavy BBHs at high redshifts will have a turnover at lower frequencies than a population of binary neutron star (BNS) or neutron star-black hole mergers. These features can be probed directly in background searches. Furthermore, if these populations have roughly comparable contributions at lower frequencies (e.g. $\sim 10-100$ Hz), then the turnover in the spectrum due to one population could distort the shape of the combined spectrum from all sources, requiring joint modeling.
@@
Furthermore, recent studies that use more realistic waveforms to make GWB predictions challenge the simple PL assumption at low frequencies because the contribution of the late inspiral phase will change the shape of the spectrum – especially the specific value of the spectral index [Renzini:2024pxt]. However, detecting such a deviation will likely require next-generation detectors. In this context, as detectors become more sensitive and resolvable CBC signals are subtracted before searching for a GWB, the residual GWB after subtraction may no longer take the same expected PL shape [Cutler:2005qq, Sachdev:2020bkk, Sharma:2020btq].
@@
Looking beyond CBC backgrounds, models like first-order phase transitions in the early Universe predict a BPL shape, where the spectral indexes before and after the break correspond to physically predictable quantities and the frequency where the spectral index changes corresponds to the temperature of the Universe when the phase transition took place [Grischuk1975, Hogan:1986dsh, Huber:2008hg, Child:2012qg, Hindmarsh:2013xza, Hindmarsh:2015qta, Hindmarsh:2017gnf, Jinno:2016vai, Weir:2017wfa, Martinovic:2020hru]. Finally, correlated magnetic noise between detectors could have a complicated spectral shape that will need to be characterized and separated from a true GWB signal [Meyers:2020qrb, janssensImpactSchumannResonances2021, janssensCorrelated11000Hz2023].
@@
Next, we turn to the merger rate vs. redshift distribution for binary black holes that contribute a GW background. Deviations from the typical Madau-Dickinson-like model could come due to lensing of high-redshift black holes [Broadhurst:2018saj, Broadhurst:2020moy, Broadhurst:2022tjm], excess star formation at high redshifts hinted at in recent James Webb Space Telescope results [finkelsteinCEERSKeyPaper2023, harikaneComprehensiveStudyGalaxies2023], or the effects of metallicity on the formation of black holes at high redshift [lehoucqAstrophysicalUncertaintiesGravitationalwave2023, perigoisStartrackPredictionsStochastic2021, kowalska-leszczynskaEffectMetallicityGravitationalwave2015, martinovicFootprintsPopulationIII2022, nakazatoGRAVITATIONALWAVEBACKGROUND2016, dvorkinMetallicityconstrainedMergerRates2016, Turbang:2023tjk].
@@
All of these added complexities, which can potentially be probed with next-generation detectors, motivate a more flexible model than a simple PL. In this paper, we use linear and spline interpolation with a variable number of knots to flexibly model the GWB spectrum and the BBH merger rate as a function of redshift. Although it may seem less informative compared to physical models used in parameter estimation, our model avoids the pitfalls of increased prior volume (and corresponding Occam penalties), reduces trials factors associated with testing multiple models, and could be used to fit correlated noise in the future as well.
@@ III Search methods overview
In this section we outline the main methods used to search for a GW background with ground-based detectors–readers already familiar with this material can skip to Sec. IV.
Common assumptions that simplify stochastic GW analyses are that the GWB is isotropic, broadband, stationary, Gaussian, and unpolarized [Allen:1999]. In this limit, a sufficient search strategy is to cross-correlate data from pairs of instruments. Although there are search strategies for detecting and characterizing a GWB that is non-Gaussian [Drasco:2002yd, smithOptimalSearchAstrophysical2018, Buscicchio:2022raf, Lawrence:2023buo], anisotropic [Allen:1996gp, Ballmer:2005uw, Mitra:2007mc, Thrane:2009fp, Ain:2018zvo, Renzini:2018vkx, Tsukada:2022nsu], narrow band [Ballmer:2005uw], or some combination of each [Thrane:2015aua, Goncharov:2018ufi, KAGRA:2021rmt, Agarwal:2023lzz, Xiao:2022uvq], we focus on the isotropic, Gaussian, unpolarized and stationary case.
A hybrid frequentist-Bayesian approach is used to perform the search [MandicThrane2012, Romano:2016dpx, Matas:2020roi, pygwb_paper]. First, an unbiased frequentist estimator for the strength of the background in individual frequency bins is constructed by cross-correlating many chunks of data [Allen:1999, Romano:2016dpx, pygwb_paper]:
$\hat{C}_{ij}(f;t)=\frac{2}{T}\frac{\textrm{Re}\left[\tilde{s}_{i}^{*}(f;t)\tilde{s}_{j}(f;t)\right]}{\Gamma_{ij}(f)S_{0}(f)},$ (3)
where $\tilde{s}_{i}(f;t)$ is the Fourier transform of the strain data in detector $i$, starting at time $t$ and ending at time $t+T$, with $T=192$ s in most recent analyses [KAGRA:2021kbb]. $\Gamma_{ij}(f)$ is the overlap reduction function [Christensen:1992wi, Flanagan:1993ix], and $S_{0}(f)=3H_{0}^{2}/(10\pi^{2}f^{3})$ is a normalization factor that makes $\langle\hat{C}_{ij}(f)\rangle=\Omega_{\textrm{GW}}(f)$ (in the absence of correlated noise [Meyers:2020qrb]).
In the limit where the GW strain amplitude is much smaller than the typical noise in the detectors [on the short, $\mathcal{O}(100\,\textrm{s})$ timescales that $\hat{C}_{ij}(f;t)$ is calculated over], the variance of $\hat{C}_{ij}(f;t)$ is given by [Romano:2016dpx]:
@@
where $\Delta f$ is the width of the frequency bins used in the analysis (0.3125 Hz in most recent LVK Collaboration results [KAGRA:2021kbb]) and $P_{i}(f;t)$ is the one-sided power spectral density (PSD) of the noise in detector $i$.
@@
In situations where the detector noise does not dominate the signal, $\sigma^{2}(f)$ [or potentially even $\sigma_{ij}^{2}(f;t)$] would need to be estimated either independently or as a part of the Bayesian step we discuss below. This is likely to be the situation for third-generation detectors, where we may no longer be in the weak signal regime. In this work, we do not account for the intermediate or large signal regime. We assume that we have an independent measurement of the detector noise PSDs. We discuss potential methods for handling the intermediate and strong signal regimes in Sec. VII.
Next, the frequentist statistic is used to perform a weighted average across frequency bins assuming a specific spectral model [Allen:1999, Romano:2016dpx, pygwb_paper], to estimate the background signal at a reference frequency. The same spectral model that fits $\hat{C}(f)$ for the spectral shape and amplitude of the background can also be used in Bayesian parameter estimation [MandicThrane2012, Callister:2017ocg, Meyers:2020qrb, pygwb_paper]. The $\hat{C}(f)$ spectrum is obtained by averaging many time segments together, and so the central limit theorem implies that it should be Gaussian distributed. Therefore, the likelihood used for a Bayesian analysis of $\hat{C}(f)$ is given by [MandicThrane2012, Romano:2016dpx, Matas:2020roi]
@@
While we discussed the broad outline of the search method in this section, for the injection studies in Sec. V we do not simulate individual time segments or time domain data–instead we calculate $\sigma^{2}(f)$ using publicly available sensitivity curves, and simulate the final $\hat{C}(f)$ using this variance.